# Breaking Data Silos: Towards Open and Scalable Mobility Foundation Models via Generative Continual Learning

### v5

@@ Abstract.
Foundation models have revolutionized fields such as natural language processing and computer vision by enabling general-purpose learning across diverse tasks and datasets. However, building analogous models for human mobility remains challenging due to the privacy-sensitive nature of mobility data and the resulting data silos across institutions. To bridge this gap, we propose MoveGCL, a scalable and privacy-preserving framework for training mobility foundation models via generative continual learning. Without sharing raw data, MoveGCL enables decentralized and progressive model evolution by replaying synthetic trajectories generated from a frozen teacher model, and reinforces knowledge retention through a tailored distillation strategy that mitigates catastrophic forgetting. To address the heterogeneity of mobility patterns, MoveGCL incorporates a Mixture-of-Experts Transformer with a mobility-aware expert routing mechanism, and employs a layer-wise progressive adaptation strategy to stabilize continual updates. Experiments on six real-world urban datasets demonstrate that MoveGCL achieves performance comparable to joint training and significantly outperforms federated learning baselines, while offering strong privacy protection. MoveGCL marks a crucial step toward unlocking foundation models for mobility, offering a practical blueprint for open, scalable, and privacy-preserving model development in the era of foundation models. To facilitate reproducibility and future research, we have released the code and models at https://github.com/tsinghua-fib-lab/MoveGCL.
Human mobility, foundation models, continual learning
journalyear: 2025copyright: ccconference: The 33rd ACM International Conference on Advances in Geographic Information Systems; November 3–6, 2025; Minneapolis, MN, USAbooktitle: The 33rd ACM International Conference on Advances inGeographic Information Systems (SIGSPATIAL ’25), November 3–6, 2025, Minneapolis, MN, USAdoi: 10.1145/3748636.3762728isbn: 979-8-4007-2086-4/2025/11ccs: Computing methodologies Machine learningccs: Information systems Mobile information processing systems11footnotetext: Equal contribution.22footnotetext: Corresponding author.
@@ 1. Introduction
In natural language processing (NLP) (kojima2022large; brown2020language; bi2024deepseek) and computer vision (CV) (esser2024scaling; liu2024sora), the rise of large pre-trained foundation models (e.g., GPT, Sora) (radford2021learning) has significantly advanced model sharing and general-purpose intelligence through centralized training and open release. However, in the critical domain of human mobility modeling (barbosa2018human; luca2021survey), the era of large foundation models has not yet arrived. This is primarily due to the highly sensitive nature of mobility trajectory data, which involves personal privacy (kong2023mobility; yuan2025learning; yuan2022activity). Such data cannot be easily shared or jointly trained across institutions due to privacy constraints and legal restrictions. As a result, mobility datasets often exist in isolated silos (yuan2025worldmove; zhu2023synmob), with different organizations and researchers relying on their own private datasets, hindering data integration, benchmarking, and collaborative model development.
@@
To address the challenge of data, recent studies have proposed different training strategies to leverage multiple datasets, and Figure 1 compares different approaches. The most common strategy is to train models separately, as shown in Figure 1(a). Recent efforts such as UniTraj (zhu2024unitraj), TrajBert (si2023trajbert), and TrajFM (lin2024trajfm) have explored joint training for unified representation and cross-city generalization (Figure 1(b)), but these models remain tightly coupled with proprietary or limited-quality datasets. TrajFM and TrajBert are typically pre-trained on restricted or private data. While UniTraj uses public data, it suffers from low semantic richness and high sparsity. Federated learning (feng2020pmf; 10598171LightTR) offers a potential solution for distributed mobility model training (Figure 1(c)), but its reliance on frequent synchronization and communication poses challenges for scalability and practical deployment. Consequently, these approaches fall short of meeting the diverse, dynamic, and multi-source demands of real-world mobility modeling, and cannot support an open ecosystem of shared models as seen in NLP and CV domains.
To truly usher in a new era of shareable and sustainable human mobility foundation models, we argue for a new collaborative paradigm. This paradigm enables multiple data holders to jointly evolve and continually build a foundation model without sharing raw data, while preserving both privacy and generalization capability. However, this vision poses several critical challenges: (1) Privacy constraints. The sensitive nature of mobility trajectory data prohibits direct data sharing. Designing a training framework that supports collaboration without exposing raw data is a fundamental yet unresolved challenge. (2) Catastrophic forgetting. Without access to past training data, the model can easily forget previously learned knowledge when updated with new mobility data, hindering long-term model evolution. (3) Data heterogeneity. Mobility data exhibits substantial variation across regions, populations, and data sources. A practical model must generalize well and dynamically adapt to such heterogeneity.
@@
  • We are the first to formalize a privacy-preserving collaborative training paradigm towards mobility foundation models, enabling decentralized model evolution without raw data sharing.
  • We propose MoveGCL, a novel framework based on generative continual learning. Its core components—knowledge distillation, mobility-aware expert routing, and layer-wise progressive adaptation—enable privacy-preserving, scalable, and adaptive trajectory model training across diverse data sources.
  • MoveGCL achieves performance comparable to joint training with full data access, and significantly outperforms federated learning baselines without data sharing.
@@ 2. Related Work
2.1. Mobility Data
2.2. Mobility Foundation Models
2.3. Continual Learning
Continual learning (kim2022theoretical; wang2024comprehensive; wang2024comprehensive), also known as lifelong learning, aims to enable models to learn from a sequence of tasks or data streams without forgetting previously acquired knowledge (yoon2021federated). A central challenge in continual learning is catastrophic forgetting (li2019learn; wickramasinghe2023continual), where the model’s performance on earlier tasks degrades as it learns new ones. Continual learning methods are typically categorized into three main types. Regularization-based methods introduce constraints on parameter updates to preserve important knowledge from earlier tasks, as seen in approaches like Elastic Weight Consolidation (EWC) and Synaptic Intelligence (SI). Replay-based methods mitigate forgetting by either storing a subset of previous data (experience replay) or generating pseudo-data (generative replay) to simulate past learning. Parameter isolation methods, on the other hand, allocate different parts of the model to different tasks, using techniques such as dynamic networks or task-specific masking to reduce interference between tasks.
@@ 3.1. Data Format
3.2. Model Training via Next-Token Prediction
@@
3.3. Training Pipeline for Mobility Foundation Model Development
@@ 4. MoveGCL
In this section, we introduce the overall framework of MoveGCL, which is shown in Figure 2. MoveGCL is built upon the paradigm of generative continual learning, as detailed in Section 4.1. We then present the core model architecture in Section 4.2, which is designed to support modular scalability and cross-city adaptability. Finally, we elaborate on the training strategy in Section 4.3, including the design of layer-wise progressive adaptation mechanism to stabilize continual updates and mitigate forgetting.
@@ 4.1. Generative Continual Learning
Generative Replay with Teacher Model.
@@
from the new dataset, where $L$ denotes its length and $\{(l^{\prime}_{i},\,t^{\prime}_{i})\}_{i=0}^{L}$ are the location–time pairs. Next, for a specific previously observed city $c_{i}$, we sample an initial location from the empirical distribution of actual locations in city $c_{i}$ conditioned on length $L$:
(4) $l_{0}\;\sim\;\rho_{\text{loc}\mid L}^{c_{i}}\,$,
where $\rho_{\text{loc}\mid L}^{c_{i}}$ denotes the empirical distribution of initial locations in previously observed city $c_{i}$ given a trajectory length $L$. We then replace $l^{\prime}_{0}$ with the sampled $l_{0}$ and generate the pseudo old‐city trajectory
@@
Distilling Knowledge to Preserve Mobility Patterns.
@@ 4.2. Model Architecture
Unified Location Encoder.
Conventional location representations often rely on discrete location IDs, which are inherently city-specific and hinder cross-city generalization. To overcome this limitation, we design a continuous location representation that embeds each location into a shared latent space, capturing transferable semantic and spatial properties across cities. This unified representation facilitates knowledge sharing and supports incremental learning across heterogeneous urban environments. Concretely, each location $l\in\mathcal{L}$ is represented by a feature vector $\mathbf{z}_{l}\in\mathbb{R}^{d}$ constructed from three key components:
(10) $\mathbf{z}_{l}=\phi_{\text{POI}}(l)\oplus\phi_{\text{lat-lon}}(l)\oplus\phi_{\text{hot}}(l),$
where $\phi_{\text{POI}}(l)\in\mathbb{R}^{d_{1}}$ denotes the Point-of-Interest (POI) embedding, capturing semantic land-use attributes such as residential, commercial, educational, or recreational functions, often indicative of mobility intent and purpose; $\phi_{\text{hot}}(l)\in\mathbb{R}^{d_{2}}$ is the mobility heat embedding, derived from public available OD flows at each location, which reflects the functional centrality of that location; $\phi_{\text{lat-lon}}(l)\in\mathbb{R}^{d_{3}}$ is the normalized latitude-longitude embedding, representing the relative spatial position of the location within the city boundary.
@@
Mixture-of-Experts Transformer.
@@
MoveGCL is built upon Mixture-of-Experts (MoE) Transformer blocks, in which each expert module is responsible for capturing specific mobility patterns. During continual learning, we introduce new experts to accommodate knowledge from new cities, and design layer-wise progressive adaptation training strategy (detail in Section 4.3). This partial parameter update strategy injects new knowledge without overwriting existing capabilities, thus alleviating catastrophic forgetting. The modularity of the MoE block also supports elastic expansion of the model as more cities are introduced.
Mobility-Aware Expert Routing.
@@
Within each layer of the MoE transformer blocks, we introduce a routing network based on a linear transformation to compute the routing weights for each expert, leveraging both the mobility descriptor feature vector $\mathbf{z}_{m}$ and the output of the self-attention submodule at that layer. The routing weights for layer $i$ are computed as:
@@
This routing strategy achieves two objectives: (1) it promotes functional specialization by directing similar motion patterns to consistent expert subsets; (2) it enables the model to discover and transfer shared mobility structures across cities, thereby enhancing generalization in multi-city scenarios. Additionally, this mobility-aware routing provides a structured inductive bias that accelerates model adaptation during incremental learning, allowing new experts to specialize rapidly with minimal interference to retained knowledge.
@@
4.3. Layer-Wise Progressive Adaptation
  • In Stage N/2, the two central layers of the model are unfrozen.
5. Results
5.1. Experimental Settings
Datasets.
We utilize human mobility datasets from multiple cities to evaluate the performance of MoveGCL. Specifically, the datasets cover over eight hundred thousand users and feature a relatively high sampling rate compared to currently available public datasets. Detailed statistics and descriptions are provided in Appendix Table 4. For each city, we randomly sample 120,000 trajectories for training, 40,000 for validation, and 40,000 for testing.
@@ 5.1. Experimental Settings
MoveGCL
0.188 0.304 0.125 0.206 0.064 0.114 0.208 0.329 0.199 0.318 0.147 0.257
MoveGCL
0.192 0.310 0.132 0.215 0.062 0.108 0.207 0.327 0.195 0.322 0.147 0.259
MoveGCL
@@
0.282 0.421 0.197 0.306 0.157 0.254 0.206 0.328 0.324 0.478 0.273 0.413
Baselines.
We compare our method with a diverse set of baselines, including traditional mobility models, federated learning-based approaches, and joint learning models with privacy-preserving mechanisms.
Parameter Settings.
The key parameters of our framework fall into three main categories.
@@ 5.2. Overall Performance
Model Atlanta Chicago Los Angeles New York Seattle Washington D.C
[…]
  • MoveGCL consistently outperforms traditional deep learning models trained independently on each dataset, demonstrating strong cross-city scalability. On average, it achieves a 8% improvement in Acc@1, highlighting its ability to generalize across diverse urban environments. This result validates the promise of mobility foundation models, which unify knowledge across cities and reduce redundancy. In contrast, training separate models for each city not only increases computational and deployment costs, but also fails to leverage shared mobility patterns across domains.
@@ 5.3. Order Invariance in Continual Learning
Figure 4. Similarity score distribution in uniqueness testing.
@@ 5.4. Privacy Evaluation
Uniqueness Testing.
Figure 4 presents the cumulative distribution of similarity scores. As shown in the figure, over 95% of the generated trajectories do not have any corresponding real trajectory with a similarity score higher than 50%. This indicates that the model’s outputs are based on the knowledge it has acquired, rather than directly copying trajectories from the training set.
Figure 5. Success rate in membership inference attack
Membership Inference Attack.
Figure 5 shows the attack results. As observed, the success rates across all datasets are approximately 50%, indicating that the classifier can hardly determine whether a trajectory was part of the training data or not based on the generated sample. These results indicate that our model is not easily susceptible to membership inference attacks.
@@
Differential Privacy.
In our experiment, we randomly select a subset of trajectories and consider two training scenarios: one in which this subset is included in the training data ($D$), and one in which it is excluded ($D^{\prime}$). For each scenario, we train $M$ on the corresponding dataset and then use each selected trajectory as a conditioning input to generate multiple synthetic trajectories. We compute a similarity score between each generated trajectory and its original conditioning trajectory. These similarity scores are modeled as two Gaussian distributions corresponding to $\mathbb{P}\bigl[M(z;\,D)\bigr]$ and $\mathbb{P}\bigl[M(z;\,D^{\prime})\bigr]$, respectively. Finally, we estimate the privacy‐budget parameters $\varepsilon$ from these distributions. As shown in Table 3, without applying any additional privacy-preserving mechanisms, MoveGCL naturally achieves a privacy budget of $\varepsilon\approx 1\text{--}2$ for 75% of randomly sampled trajectories. This level is generally considered an acceptable operating point for generative models (lin2020using); for example, Apple adopts a privacy budget of $\varepsilon=4.0$ ***https://www.apple.com/privacy/docs/Differential_Privacy_Overview.pdf.

### v6

@@ Abstract.
Human mobility is a fundamental pillar of urban science and sustainability, providing critical insights into energy consumption, carbon emissions, and public health. However, the discovery of universal mobility laws is currently hindered by the “data silo” problem, where institutional boundaries and privacy regulations fragment the necessary large-scale datasets. In this paper, we propose MoveGCL, a transformative framework that facilitates collaborative and decentralized mobility science via generative continual learning. MoveGCL enables a distributed ecosystem of data holders to jointly evolve a foundation model without compromising individual privacy. The core of MoveGCL lies in its ability to replay synthetic trajectories derived from a generative teacher and utilize a mobility-pattern-aware Mixture-of-Experts (MoE) architecture. This allows the model to encapsulate the unique characteristics of diverse urban structures while mitigating the risk of knowledge erosion (catastrophic forgetting). With a specialized layer-wise progressive adaptation strategy, MoveGCL ensures stable convergence during the continuous integration of new urban domains. Our experiments on six global urban datasets demonstrate that MoveGCL achieves performance parity with joint training, a previously unattainable feat under siloed conditions. This work provides a scalable, privacy-preserving pathway toward Open Mobility Science, empowering researchers to address global sustainability challenges through cross-institutional AI collaboration. To facilitate reproducibility and future research, we have released the code and models at https://github.com/tsinghua-fib-lab/MoveGCL.
11footnotetext: Equal contribution.
@@ 1. Introduction
The study of human mobility is more than a technical task; it is a cornerstone of urban systems theory and sustainability science (xu2025using; huang2015defining). Understanding how billions of people move through urban environments is essential for optimizing transportation efficiency, reducing the global carbon footprint, and managing epidemiological risks (nouvellet2021reduction; zeng2024estimating; zheng2023spatial). In the era of Foundation Models, fields such as natural language processing (NLP) (kojima2022large; brown2020language; bi2024deepseek) and computer vision (CV) (esser2024scaling; liu2024sora) have undergone a paradigm shift. This evolution presents an unprecedented opportunity to move beyond insular empirical studies, which often focus on isolated datasets and specific cities, toward a unified, generalizable science of mobility.
However, a critical barrier remains: the fragmentation of mobility data across organizational and national boundaries (yuan2025worldmove; zhu2023synmob). Unlike the public-domain text used in NLP and CV, mobility data is inherently sensitive, held by disparate telecommunications operators or transit agencies in isolated “data silo” due to privacy concerns and stringent regulatory frameworks (e.g., GDPR) (kong2023mobility; yuan2025learning). This isolation prevents the scientific community from training large-scale models that can generalize across different urban morphologies and socio-economic contexts. Consequently, the field is stuck in a paradigm where models are trained on small, localized datasets, failing to capture the complex, universal patterns of human movement.
@@
To address these data barriers, recent studies have proposed different training strategies to leverage multiple datasets, and Figure 1 compares different approaches. The most common strategy is to train models separately, as shown in Figure 1(a). Recent efforts such as UniTraj (zhu2024unitraj), TrajBert (si2023trajbert), and TrajFM (lin2024trajfm) have explored joint training for unified representation and cross-city generalization (Figure 1(b)), but these models remain tightly coupled with proprietary or limited-quality datasets. TrajFM and TrajBert are typically pre-trained on restricted or private data. While UniTraj uses public data, it suffers from low semantic richness and high sparsity. Federated learning (feng2020pmf; 10598171LightTR) offers a potential solution for distributed mobility model training (Figure 1(c)), but its reliance on frequent synchronization and communication poses challenges for scalability and practical deployment. Consequently, these approaches fall short of meeting the diverse, dynamic, and multi-source demands of real-world mobility modeling, and cannot support an open ecosystem of shared models as seen in NLP and CV domains.
To truly usher in an era of shareable and sustainable human mobility science, we argue for a new collaborative paradigm. This paradigm enables multiple data holders to jointly evolve and continually build a foundation model without sharing raw data, while preserving both privacy and generalization capability. However, this vision poses several critical challenges: Privacy preservation: How to achieve cross-institutional collaboration without any raw data exchange. (2) Cumulative knowledge retention: How to prevent “catastrophic forgetting” so that the model can accumulate knowledge from diverse urban systems over time. (3) Scientific heterogeneity: How to design an architecture that captures the unique characteristics of different urban structures (e.g., the grid layouts of New York vs. the radial patterns of Paris).
@@
  • We formalize the first privacy-preserving collaborative training paradigm for mobility foundation models, enabling the scientific community to build unified models across data silos.
  • We propose MoveGCL, a novel framework based on generative continual learning. Its core components—knowledge distillation, mobility-aware expert routing, and layer-wise progressive adaptation address catastrophic forgetting and urban heterogeneity, facilitating the long-term accumulation of mobility knowledge.
  • We demonstrate on six diverse global urban datasets that MoveGCL achieves performance comparable to centralized joint training, providing a practical pathway toward Open Mobility Science.
@@ 2. Related Work
Due to space limitations, we provide a concise review here and defer the extended discussion to Appendix B.
Mobility Data
Mobility data is commonly represented as aggregated flows or individual trajectories (barbosa2018human; yuan2025learning; zhang2025noise). While flow data is relatively accessible and widely used in urban analytics (rong2024interdisciplinary; zhang2017deep), trajectory data is often fragmented due to privacy constraints and institutional silos (kong2023mobility; yuan2025learning). Existing real-world and open trajectory datasets are typically limited in city coverage, duration, or sampling density, and even recent global-scale datasets may suffer from inconsistent resolution and quality. Generative AI has enabled synthetic mobility datasets (zhu2023synmob; yabe2024yjmob100k; yuan2025worldmove), but matching real trajectories remains challenging, especially in behavioral diversity and temporal continuity. As a result, many studies rely on restricted-access data and cannot release the trajectories they use (schlapfer2021universal).
Mobility Foundation Models
Recent work has begun to build foundation models for urban and mobility applications (yuan2024urbandit; chai2025mobiworld; han2025trajmoe; zhou2024urban; zhang2024urban; choudhury2024towards). Most efforts focus on aggregated mobility signals and demonstrate strong cross-city transfer via unified spatio-temporal pre-training (li2024urbangpt; yuan2024uniflow; yuan2024unist; li2024opencity). In contrast, foundation models for individual mobility remain less mature due to limited representative trajectory corpora and the complexity of human behavior (yuan2025learning; zhang2025noise; long2024universal). LLM-based trajectory generation has also been explored (shao2024beyond; gong2024mobility; ju2025trajllm; feng2024agentmove; jiawei2024large), but the modality gap suggests the need for native mobility foundation models that can later be aligned with language-based reasoning.
Continual Learning
Continual (lifelong) learning aims to update models on sequential data without degrading previously learned capabilities (kim2022theoretical; wang2024comprehensive; yoon2021federated). A key challenge is catastrophic forgetting (li2019learn; wickramasinghe2023continual). Existing methods typically fall into regularization, replay, and parameter-isolation families, which trade off memory, computation, and stability when adapting to new data streams.
@@ 3. Preliminaries
Model Training via Next-Token Prediction
@@
Training Pipeline for Mobility Foundation Model Development
@@ 4. MoveGCL
In this section, we introduce the overall framework of MoveGCL, which is shown in Figure 2.
@@ 4.1. Generative Continual Learning
Generative Replay with Teacher Model
@@
from the new dataset, where $L$ denotes its length and $\{(l^{\prime}_{i},\,t^{\prime}_{i})\}_{i=0}^{L}$ are the location–time pairs. Next, for a specific previously observed city $c_{i}$, we sample an initial location from the empirical distribution of actual locations in city $c_{i}$ conditioned on length $L$: $l_{0}\;\sim\;\rho_{\text{loc}\mid L}^{c_{i}}\,$, where $\rho_{\text{loc}\mid L}^{c_{i}}$ denotes the empirical distribution of initial locations in previously observed city $c_{i}$ given a trajectory length $L$. We then replace $l^{\prime}_{0}$ with the sampled $l_{0}$ and generate the pseudo old‐city trajectory
@@
Distilling Knowledge to Preserve Mobility Patterns
@@ 4.2. Model Architecture
Unified Location Encoder
Conventional location representations often rely on discrete location IDs, which are inherently city-specific and hinder cross-city generalization. To overcome this limitation, we design a continuous location representation that embeds each location into a shared latent space, capturing transferable semantic and spatial properties across cities. This unified representation facilitates knowledge sharing and supports incremental learning across heterogeneous urban environments. Concretely, each location $l\in\mathcal{L}$ is represented by a feature vector $\mathbf{z}_{l}\in\mathbb{R}^{d}$ constructed from three key components: $\mathbf{z}_{l}=\phi_{\text{POI}}(l)\oplus\phi_{\text{lat-lon}}(l)\oplus\phi_{\text{hot}}(l),$ where $\phi_{\text{POI}}(l)\in\mathbb{R}^{d_{1}}$ denotes the Point-of-Interest (POI) embedding, capturing semantic land-use attributes such as residential, commercial, educational, or recreational functions, often indicative of mobility intent and purpose; $\phi_{\text{hot}}(l)\in\mathbb{R}^{d_{2}}$ is the mobility heat embedding, derived from public available OD flows at each location, which reflects the functional centrality of that location; $\phi_{\text{lat-lon}}(l)\in\mathbb{R}^{d_{3}}$ is the normalized latitude-longitude embedding, representing the relative spatial position of the location within the city boundary.
@@
Mixture-of-Experts Transformer
@@
MoveGCL is built upon Mixture-of-Experts (MoE) Transformer blocks, in which each expert module is responsible for capturing specific mobility patterns. During continual learning, we introduce new experts to accommodate knowledge from new cities, and design layer-wise progressive adaptation training strategy (detail in Section 4.3). This partial parameter update strategy injects new knowledge without overwriting existing capabilities, thus alleviating catastrophic forgetting. The modularity of the MoE block also supports elastic model expansion as more cities are introduced.
Mobility-Aware Expert Routing
@@
Within each layer of the MoE blocks, we introduce a routing network based on a linear transformation to compute the routing weights for each expert, leveraging both the mobility descriptor feature vector $\mathbf{z}_{m}$ and the output of the self-attention submodule at that layer. The routing weights for layer $i$ are computed as:
@@
This routing strategy achieves two goals: (1) it promotes functional specialization by directing similar motion patterns to consistent expert subsets; (2) it enables the model to discover and transfer shared mobility structures across cities, thus enhancing generalization in multi-city scenarios. Additionally, this mobility-aware routing provides a structured inductive bias that accelerates model adaptation during incremental learning, allowing new experts to specialize rapidly with minimal interference to retained knowledge.
@@
MoveGCL (FullTune) 0.188 0.304 0.125 0.206 0.064 0.114 0.208 0.329 0.199 0.318 0.147 0.257
MoveGCL (ExpertTune) 0.192 0.310 0.132 0.215 0.062 0.108 0.207 0.327 0.195 0.322 0.147 0.259
@@
4.3. Layer-Wise Progressive Adaptation
  • In Stage N/2, the two central layers of the model are unfrozen.
5. Results
5.1. Experimental Settings
Datasets. We utilize human mobility datasets from multiple cities to evaluate the performance of MoveGCL. Specifically, the datasets cover over eight hundred thousand users and feature a relatively high sampling rate compared to currently available public datasets. Detailed statistics and descriptions are provided in Appendix Table 3. For each city, we randomly sample 120,000 trajectories for training, 40,000 for validation, and 40,000 for testing.
Baselines and Parameter Settings. We compare our method with a diverse set of baselines, including traditional mobility models, federated learning-based approaches, and joint learning models with privacy-preserving mechanisms. Appendix A.2 provides detailed descriptions of the baselines, and Appendix A.3 summarizes the parameter settings.
@@ 5.2. Overall Performance
  • MoveGCL consistently outperforms traditional deep learning models trained independently on each dataset, demonstrating strong cross-city scalability. On average, it achieves an 8% improvement in Acc@1, highlighting its ability to generalize across diverse urban environments. This result validates the promise of mobility foundation models, which unify knowledge across cities and reduce redundancy. In contrast, training separate models for each city not only increases computational and deployment costs, but also fails to leverage shared mobility patterns across domains.
@@ 5.3. Order Invariance in Continual Learning
To assess the robustness of MoveGCL in real-world deployment scenarios, we investigate its sensitivity to the order in which data from different cities is introduced during continual learning. As shown in Appendix Table 4, the performance of MoveGCL remains remarkably consistent, with the vast majority of metrics showing deviations of less than 5% across original and reversed sequences. This empirical finding confirms the order-invariant learning behavior of our model, demonstrating that MoveGCL can integrate new city data without disrupting previously acquired knowledge, even when the order of exposure varies significantly. We refer the reader to the detailed analysis in Appendix C.1.
@@ 5.4. Privacy Evaluation
Figure 4. Similarity score distribution in uniqueness testing.
Figure 5. Success rate in membership inference attack
@@
Due to space limitations, we summarize the key conclusions here, and provide detailed experimental settings for the privacy evaluation in Appendix C.2.
Uniqueness Testing. Figure 4 presents the cumulative distribution of similarity scores. As shown in the figure, over 95% of the generated trajectories do not have any corresponding real trajectory with a similarity score higher than 50%. This indicates that the model’s outputs are based on the knowledge it has acquired rather than directly copying trajectories from the training set.
Membership Inference Attack. Figure 5 shows the attack results. As observed, the success rates across all datasets are approximately 50%, indicating that the classifier can hardly determine whether a trajectory was part of the training data or not based on the generated sample. These results indicate that our model is not easily susceptible to membership inference attacks.
@@
Differential Privacy. As shown in Table 2, without applying any additional privacy-preserving mechanisms, MoveGCL naturally achieves a privacy budget of $\varepsilon\approx 1\text{--}2$ for 75% of randomly sampled trajectories. This level is generally considered an acceptable operating point for generative models (lin2020using); for example, Apple adopts a privacy budget of $\varepsilon=4.0$ ***https://www.apple.com/privacy/docs/Differential_Privacy_Overview.pdf.